OpenStudy (anonymous):

If 3-4y=5x-6, then y=? ...with steps please :)

OpenStudy (saifoo.khan):

4y = -5x +9

OpenStudy (saifoo.khan):

y = -5x + 9 --- --- 9 9

OpenStudy (anonymous):

well first subtract 3 from each side to give you -4y=5x-9 now divide each side by -4 for y=-5/4x+9/4

OpenStudy (saifoo.khan):

y = -5x + 1 --- 9

OpenStudy (anonymous):

The answer the book gave is -5x-9/4

OpenStudy (anonymous):

y =- (5x - 9)/4

OpenStudy (anonymous):

saifoo went a bit wrong there...

OpenStudy (anonymous):

Estudier can you show me the steps to your answer.

OpenStudy (anonymous):

cough cough look at elizateds answer cough cough

OpenStudy (anonymous):

3-4y=5x-6 -4y = 5x -9 y =-(5x-9)/4 (I put the minus at the top, makes no difference)

OpenStudy (anonymous):

@elizated Yes u have it right as well, however it is customary to add fractions with the same denominator.